OpenStudy (anonymous):

Motion Problems - D = R x T Four pencils and five pens cost $2.00. Three pencils and four pens cost $1.58. Find the coast of the pencil and the cost of the pen.

OpenStudy (anonymous):

Let a = cost of pencil Let b = cost of pen 4a + 5b = 2 3a + 4b = 1.58 Now we have two equations and two unknow.

OpenStudy (anonymous):

Can u solve it now

OpenStudy (anonymous):

I still need help

OpenStudy (anonymous):

We can solve this, by using elimination Lets call this equation 1) 4a + 5b = 2 Lets call this equation 2) 3a + 4b = 1.58 Multipy equation 1 by (3) Multipy equation 2 by (-4) this gives us, 12a + 15b = 6 -12a - 16b = -6.32 we can now add the two equations together to eliminate the (a), doing so leaves us with -b = -0.32 divide both sides by -1, to get b = $0.32 We can now plug (b) into equation 1 or 2, to get (a) I will go with (1) 4a + 5(0.32) = 2 4a + 1.6 = 2 subtract 1,6 from both sides 4a = 0.4 divide both sides by 4 a = $0.1

OpenStudy (anonymous):

The cost for pencil = $0.1. The cost for pen = $0.32